TC= 103+15q-6q^2+q^3




Find VC , MC, AVC




Value of FC + VC when out put (q )= 0.5

1
Expert's answer
2022-02-15T16:13:00-0500

VC = "15q-6{q^2}+{q^3}"


MC = "\frac{dMC}{dq} =15-12q+3{q^2}"


AVC = "\frac{VC}{q}"


AVC = "\frac{15q-6{q^2}+{q^3}}{q}"


AVC = "15-6q+{q^2}"

Value of FC + VC when out put (q )= 0.5

FC = 103

VC = "15q-6{q^2}+{q^3}"

VC = "15(0.5)-6({0.5^2})+{0.5^3}"

VC = 6.125

FC + VC = 103+6.125

Value = 109.125